OpenStudy (anonymous):

Deion pitches a baseball at a speed of 48 miles per hour. The distance between Deion and the batter is 60 feet. Based on the speed of the pitch, approximately how much time does it take the ball to travel between Deion and the batter? 1 mile = 5,280 feet between 0 and 0.5 seconds between 0.5 seconds and 1 second between 1 second and 1.5 seconds between 1.5 seconds and 2 seconds Kahlina is traveling to India. She estimates that she will need to have 20,000 Indian rupees with her when she arrives in India. She thinks that she can use the conversion rates below.

OpenStudy (anonymous):

Rate A: 1 US dollar is equal to 61.25 Indian rupees Rate B: 1 Indian rupee is equal to 0.0163 US dollars She uses both conversion rates to attempt to determine how many US dollars she needs to convert before leaving. Expression A: Expression B: She determines that one of the conversion rates must be incorrect because the two expressions produce different results. Which best describes her conclusion? Her conclusion is correct, the conversion rates are not equivalent. Her expressions are correct based on the incorrect rates. Her conclusion is correct, the conversion rates are not equivalent. However, one or both of her expressions are incorrect. Her conclusion is incorrect, the conversion rates are equivalent. Expression A incorrectly converts rupees to dollars. Her conclusion is incorrect, the conversion rates are equivalent. Expression B incorrectly converts rupees to dollars.

OpenStudy (anonymous):

\[\frac{ 1hour }{ 48 miles}*\frac{60 minutes }{ 1hour }*\frac{ 60seconds }{ 1minute }*\frac{ 1mile }{ 5280feet}*\frac{ 60feet }{ 1 }= answer\]